# AceCoder: Acing Coder RL via Automated Test-Case Synthesis

### v3

@@ 4.5 Ablation Studies
We also conduct experiments to investigate how filtering the test cases with a proxy model can affect the results. As shown in  Table 5, training RM on data after the filtering improve the performance significantly, especially for those hard code questions like MBPP-Plus and BigCodeBench-Hard (C/I). We believe this is because the test case filtering can ensure the remaining ones are consistent with each other and thus point to the same implicit program, which improves the quality of the rewards.
@@
Inspired by the recent DeepSeek-R1 (Guo et al., 2025), we also conduct the RL directly from the base model without any SFT. It turns out we get huge improvements when using rule-based rewards. For example, we get $25.0$ points of improvements on HumanEval-Plus after training only 6 hours from the Base Model, which is way more efficient that the large-scale SFT. What’s more, the AceCoder Rule improve the BigCodeBench-Instruct-Full’s performance from $40.2$ to $43.2$, nearly the same performance with DeepSeek-R1-Distill-Qwen-32B ($43.9$) which was directly distilled from the DeepSeek-R1 Model. This further consolidates the finding of DeepSeek-Zero. However, we do find that using reward models for RL tuning can lead to worse results. We attribute this to the potential reward hacking during the tuning.
@@ 5.4 Reinforcement Learning for LLM
Given that coding is an inherently verifiable task, recent studies have explored RL techniques that leverage direct execution accuracy as a reward signal. PPOCoder (Shojaee et al., 2023b) and CodeRL (Le et al., 2022) demonstrated the effectiveness of PPO-based RL for coding tasks, while RLEF (Gehring et al., 2024) extended this approach to multi-turn settings by incorporating execution feedback at each step. StepCoder (Dou et al., 2024a) refined the reward mechanism by assigning rewards at a more granular level, considering only successfully executed lines of code. Additionally, DSTC (Liu et al., 2024d) explored the application of Direct Preference Optimization (DPO) to code generation by using self-generated test cases and programs.
@@ 6 Conclusion
We are first work to automate large-scale test-case synthesis and adopt them to train coder language models. Without relying on the most advanced model, our data collection pipeline can still produce very high-quality verifiable code data, which empowers the training of reward model and coder model through reinforcement learning. Though our work demonstrates huge improvement in Best-of-N experiments, the improvement on RL training is less prominent. We believe future work should further our reward model training to improve its robustness to further the results.
@@ Appendix A Appendix
A.1 Prompt Template

### v4

@@ 3.5 Ablation Studies
We also conduct experiments to investigate how filtering the test cases with a proxy model can affect the results. As shown in  Table 7, training RM on data after the filtering improves the performance significantly, especially for those hard code questions like MBPP-Plus and BigCodeBench-Hard (C/I). We believe this is because the test case filtering can ensure the remaining ones are consistent with each other and thus point to the same implicit program, which improves the quality of the rewards.
@@
Inspired by the recent DeepSeek-R1 (Guo et al., 2025), we also conduct the RL directly from the base model without any SFT. It turns out we get huge improvements when using rule-based rewards. For example, we get $25.0$ points of improvements on HumanEval-Plus after training only 6 hours from the Base Model, which is way more efficient than large-scale SFT. What’s more, the AceCoder Rule improve the BigCodeBench-Instruct-Full’s performance from $40.2$ to $43.2$, nearly the same performance with DeepSeek-R1-Distill-Qwen-32B ($43.9$) which was directly distilled from the DeepSeek-R1 Model. This further consolidates the finding of DeepSeek-Zero. However, we do find that using reward models for RL tuning can lead to worse results. We attribute this to the potential reward hacking during the tuning process.
@@ 4.2 Reinforcement Learning for LLM
Given that coding is an inherently verifiable task, recent studies have explored RL techniques that leverage direct execution accuracy as a reward signal. PPOCoder (Shojaee et al., 2023b) and CodeRL (Le et al., 2022) demonstrated the effectiveness of PPO-based RL for coding tasks, while RLEF (Gehring et al., 2024) extended this approach to multi-turn settings by incorporating execution feedback at each step. StepCoder (Dou et al., 2024a) refined the reward mechanism by assigning rewards at a more granular level, considering only successfully executed lines of code. Additionally, DSTC (Liu et al., 2024e) explored the application of Direct Preference Optimization (DPO) to code generation by using self-generated test cases and programs.
@@ 5 Conclusion
We introduced AceCoder as the first approach to reward model training and RL tuning for code generation using large-scale, reliable test case synthesis. Our data pipeline produces high-quality verifiable code without relying on the most advanced models, enabling effective reward model training and reinforcement learning. Our method significantly improves Best-of-N performance. However, RL training gains are less pronounced, leaving it as a future work to enhance.
Limitations
Test Case Synthesis
Despite our efforts to enhance the reliability of synthesized test cases through prompt engineering and filtering with a reference solution, inaccuracies can still arise. These errors may stem from an incorrect reference solution or test cases that are too simple, failing to capture challenging edge cases. Consequently, passing all test cases does not necessarily guarantee a program’s correctness, leading to noise in the reward model training and reinforcement learning (RL) tuning signals. To address this, future work can leverage stronger large language models (LLMs) to synthesize more rigorous test cases, ensuring the inclusion of harder corner cases. Additionally, using more advanced coding LLMs to generate reference solutions could further improve test case filtering, preserving only high-quality examples.
Reinforcement Learning for Coding
In this paper, we explored RL tuning using three models and two types of rewards: RM-based and rule-based. While significant improvements are observed when tuning Qwen2.5-7B-Instruct and Qwen2.5-Coder-7B-Base, tuning on Qwen2.5-Coder-7B-Instruct exhibited less pronounced gains due to its strong ability originally. This suggests that the current reward signals may still contain noise. Furthermore, there remains considerable room for improvement, particularly in tuning the Qwen2.5-Coder-7B-Base. Given recent advancements in models such as DeepSeek-R1, future work could further refine RL tuning strategies to achieve better performance with a more fine-grained reward design.
Ethical Statements
This work fully complies with the ACL Ethics Policy. We declare that there are no ethical issues in this paper, to the best of our knowledge.
@@ Appendix A Appendix
A.1 More related works
LLM for Code Generation
Reward Models
A.2 AceCode-RM Model Breakdown
In this section, we present the different models involved in the training and evaluation process of AceCode-RM in a concise table for further clarification.
AceCodeRM-7B
Backbone Qwen2.5-Coder-7B-Instruct
Training Responses (Preference Pairs) Generation Qwen2.5-Coder-7B-Instruct
Inference Model Any Model (Qwen2.5-Coder-7B-Instruct, Llama-3.1-8B-Instruct, etc.)
AceCodeRM-32B
Backbone Qwen2.5-Coder-32B-Instruct
Training Responses (Preference Pairs) Generation Qwen2.5-Coder-32B-Instruct
Inference Model Any Model (Qwen2.5-Coder-7B-Instruct, Llama-3.1-8B-Instruct, etc.)
Table 9: Models used during the training and evaluation process of AceCode-RM.
A.3 Prompt Template
@@
A.4 Case Studies on HumanEval
Figure 2: In this example, the RL model took a more advanced approach and attempted to use regular expression matching. However, the regular expression it came up is not correct as it did not include whitespace and did not handle the constraint "there should not be more than three digits (’0’-’9’) in the file’s name" correctly.
Figure 3: In this example, the RL model correctly caught the error where the baseline did not consider the case where the whole string has to be repeated in order to create a palindrome.
Figure 4: In this example, while both codes would have identical output, the baseline’s output is slow due to the recursive calls. By using a for loop instead, the RL model’s code’s runtime is half of that of the baseline’s. Therefore, it passed the test whereas the baseline’s code got a time-out.